OpenStudy (anonymous):

solve each system of equations by the substitution method. Tell whether each system is inconsistent or consistent. If the system is consistent, tell whether it is independent or dependent x+3y=18 3x-y=-6

OpenStudy (bahrom7893):

i only see one system.

OpenStudy (anonymous):

x=0 i know that much but idk bout y

OpenStudy (bahrom7893):

if x=0, y = 6 just plug in 0 into any of the equations to get y.

OpenStudy (anonymous):

huh

OpenStudy (bahrom7893):

0+3y=18 3y=18 y=18/3=6

OpenStudy (bahrom7893):

or 0-y=-6 -y=-6 y=6